# Mechanistic Interpretability of Emotion Inference in Large Language Models

### v1

@@ 8 Emotion-Appraisal Mappings
Figure 6 shows the similarity score of emotion vectors with two appraisal vectors, i.e. the pleasantness and other-agency, throughout the layers of Llama 3.2 1B. Notably, we observe psychologically plausible appraisal-emotion mappings across all layers. However, the projection strength peaks in the early layers and fades to near zero in the final layer, suggesting orthogonality in the final layers.
We hypothesize that in the earlier layers, there exists a meaningful structure on appraisal and emotion concepts, which aligns with our expectations from the appraisal theory. However, these concepts gradually decouple as the processing progresses through the network, and by the final layers, they become orthogonal and fully decoupled, reflecting the specialization of the network toward higher-level tasks. We finish this section by drawing the connection to our findings in previous sections. Notice that the decoupling starts around the critical layer, e.g. layer 10 in Llama 3.2 1B, which we identified in previous sections. Therefore, we conclude that the appraisals build a foundation to understand emotion representations in LLM hidden states, but the structure vanishes as we progress through the network.
@@ 9 Intervention on Appraisal Concepts
After finding the appraisal vectors, we investigate the possibility of indirectly modifying the emotion of an input example by modulating its appraisals within the model representations. For this purpose, we need to isolate the role of each appraisal $a$, by considering its associated latent vector $\mathbf{v}_{a}$ and distinguishing it from other appraisal vectors. More precisely, we define $\mathbf{V}_{-a}:=\left[\mathbf{v}_{1},\dots,\mathbf{v}_{a-1},\mathbf{v}_{a+1},\dots,\mathbf{v}_{n}\right]$ by contacting all appraisal vectors except $\mathbf{v}_{a}$. Next, we introduce the unique effect vector of appraisal $a$ as $\mathbf{z}_{a}:=(\mathbf{I}-\mathbf{P}_{-a})\mathbf{v}_{a}$, where $\mathbf{P}_{-a}=\mathbf{V}_{-a}(\mathbf{V}^{\top}_{-a}\mathbf{V}_{-a})^{-1}\mathbf{V}^{\top}_{-a}$ is the projection matrix onto the column space of $\mathbf{V}_{-a}$ and $\mathbf{I}\in\mathbb{R}^{d\times d}$ is the identity matrix. We perform appraisal modulation by injecting $\mathbf{z}_{a}$ into the model’s latent representation. More formally, the intervention is expressed as
@@
where $\mathbf{x}$ on the RHS is the original latent representation, e.g., a hidden state vector from a specific layer and $\beta$ is a scaling factor controlling the strength of the concept modulation. Notice that a positive $\beta$ corresponds to an appraisal promotion while a negative $\beta$ has the opposite effect of appraisal demotion. To measure the success of interventions, we evaluate the new emotion label derived by this modification and repeat this procedure across all examples. This modification is applicable to each layer of the model.
Figure 7 illustrates concept modulation results with different magnitudes of $\beta$ on layer $9$ of Llama 3.2 1B. We observe a remarkable alignment with theoretical and intuitive expectations. For instance, we observe that increasing the pleasantness appraisal promotes both joy and pride, aligning with the fact that both of these emotions have high associations with the appraisal.
@@ 10 Discussion
A key distinction of our work is that we grounded it on psychological theory and applied MI analysis on in-the-wild examples, rather than relying on synthetically generated simplistic structures, as seen in prior studies Merullo et al. (2024). For example, Wang et al. (2023a) study the Indirect Object Identification task, by considering a fixed input structure, such as “person1 and person2 had fun at school. person2 gave a ring to” where the model is expected to predict “person1”. Hanna et al. (2023) study the “Greater-than" task using a dataset of examples like “The war lasted from 1517 to 15”, where the model is expected to predict any two-digit number larger than 17. While it is possible to identify specialized circuits for such tasks, interpreting them effectively requires structured inputs, making it challenging to generalize findings to more naturalistic settings
Beyond these contributions, our work also highlights new opportunities for future research. Despite significant advancements in understanding human emotions, debates persist regarding the definition of emotion, the role of cognition in emotion, and the mechanisms underlying emotion inference Ortony et al. (2022); Ellsworth and Scherer (2003); Moors (2013); Barrett (2017). In parallel, cognitive neuroscience has explored the neural basis of emotion in support of differing theoretical perspectives Kragel et al. (2024). The study of LLMs, combined with insights from emotion theory and neuroscience, opens a unique intersection for advancing our understanding of emotions Sievers and Thornton (2024).
@@ 11 Limitations
In this study, we build upon the linear representation hypothesis Mikolov et al. (2013a, b); Levy and Goldberg (2014); Elhage et al. (2022)—the idea that high-level concepts are encoded linearly within model representations Park et al. (2024). This hypothesis is particularly appealing because, if true, it could enable simple and effective methods for interpreting and controlling LLMs—an approach we leveraged to localize and manipulate latent emotion representations. However, despite recent notable efforts to formalize the notions of linearity Park et al. (2024) and orthogonality Jiang et al. (2024b) in model representations, further research is needed to enhance clarity and robustness in this area.
@@ 12 Ethical Impact Statement
This study re-analyzes previously collected, de-identified data that had already undergone ethical review. The dataset is used for investigating the inner mechanisms by which auto-regressive LLMs process emotion. However, caution must be exercised when generalizing these results to models not examined in this work, to superficially similar tasks, or to different languages. Our analysis highlights potential concerns for those deploying LLMs in high-stakes affective domains or for generating emotionally charged content. Given the risks associated with emotional manipulation by LLMs, it is crucial to develop a deeper understanding of how these models process emotions. To this end, we advocate for further research in this domain to ensure that LLMs align with ethical standards and human-centered AI principles.
@@ A.1 Dataset Details
For this project, we employed the crowd-enVENT dataset. Troiano et al. (2023) developed the dataset by asking crowdsource writers to share an event that made them feel a particular emotion. The participants were then asked to evaluate their subjective experiences during that event, including their perceived appraisals. Both were rated on a 5-point Likert scale with 5 representing the highest agreement.
@@ A.3 Task Details
For the rest of our analysis, we only focus on the correctly classified examples, which ensures at least 2,700 data points or more across different model architectures. We apply this filtering method to ensure that the samples selected for experimentation are ones where the LLM understood the emotion labeling task, allowing us to properly investigate the underlying mechanisms that led to the models’ selected emotion label. However, we acknowledge that some emotion classes have a limited number of examples after filtering, which may present constraints in our experiments.
@@ Appendix F Further Details on Appraisal Modulation
On the other hand, We observe a remarkable alignment with theoretical and intuitive expectations in the distribution shifts associated with interventions on middle layers (specifically layers 9-11). For instance, we observe that increasing the pleasantness appraisal promotes both joy and pride, aligning with the fact that both of these emotions have high associations with the appraisal. Also see 35 for a different visualization.
@@
These findings provide strong evidence that the mid-layers in Llama 3.2 1B directly contribute to cognitive processes related to emotions.
@@ Appendix G Code and Compute Resources
Our experiments are conducted using GPU-accelerated compute resources, with hardware such as NVIDIA A100 GPUs. For larger models, our studies are feasible on GPUs with at least 40GB of VRAM, with the full experiment running in approximately 24 hours. For smaller models, GPUs with 12GB of VRAM are sufficient to carry out our analyses efficiently. Our implementation and experiment code are publicly available at: Emo-LLM Github Repository.

### v2

@@ 8 Emotion-Appraisal Mappings
Figure 6 shows the similarity score of emotion vectors with two appraisal vectors, i.e. the pleasantness and other-agency, throughout Llama 3.2 1B layers. Notably, we observe psychologically plausible appraisal-emotion mappings across all layers. However, the projection strength peaks in the early layers and fades to near zero in the final layer, suggesting orthogonality in the final layers.
We hypothesize that in the earlier layers, there exists a meaningful structure capturing appraisal and emotion concepts, which aligns with our expectations from the appraisal theory. However, these concepts gradually decouple as the processing progresses through the network, and by the final layers, they become orthogonal and fully decoupled, reflecting the specialization of the network toward higher-level tasks. We finish this section by drawing the connection to our findings in previous sections. Notice that the decoupling starts around the critical layers, e.g. layers 9 and 10 in Llama 3.2 1B, which we identified in previous sections. Therefore, our observations suggest that the appraisals build a foundation to understand emotion representations in LLM hidden states, but the structure vanishes as we progress through the network.
@@ 9 Intervention on Appraisal Concepts
After finding the appraisal vectors, we investigate the possibility of indirectly modifying the emotion of an input example by modulating its appraisals within the model representations. For this purpose, we need to isolate the role of each appraisal $a$, by considering its associated latent vector $\mathbf{v}_{a}$ and distinguishing it from other appraisal vectors. More precisely, we define ${\mathbf{V}_{-a}:=\left[\mathbf{v}_{1},\dots,\mathbf{v}_{a-1},\mathbf{v}_{a+1},\dots,\mathbf{v}_{n}\right]}$ by concatenating all appraisal vectors except $\mathbf{v}_{a}$. Next, we introduce the unique effect vector of appraisal $a$ as ${\mathbf{z}_{a}:=(\mathbf{I}-\mathbf{P}_{-a})\mathbf{v}_{a}}$, where ${\mathbf{P}_{-a}=\mathbf{V}_{-a}(\mathbf{V}^{\top}_{-a}\mathbf{V}_{-a})^{-1}\mathbf{V}^{\top}_{-a}}$ is the projection matrix onto the column space of $\mathbf{V}_{-a}$ and ${\mathbf{I}\in\mathbb{R}^{d\times d}}$ is the identity matrix. We perform appraisal modulation by injecting $\mathbf{z}_{a}$ into the model’s latent representation. More formally, the intervention is expressed as:
@@
where $\mathbf{x}$ on the RHS is the original latent representation, e.g., a hidden state vector from a specific layer, and $\beta$ is a scaling factor controlling the strength of the concept modulation. Notice that a positive $\beta$ corresponds to an appraisal promotion while a negative $\beta$ has the opposite effect of appraisal demotion. To measure the success of interventions, we evaluate the new emotion label derived from this modification and repeat this procedure across all examples. Notice that, this modification is applicable to any layer of the model.
Figure 7 illustrates concept modulation results with different magnitudes of $\beta$ on layer $9$ of Llama 3.2 1B. Notably, we observe a remarkable alignment with theoretical and intuitive expectations. For instance, increasing the pleasantness appraisal promotes both joy and pride, aligning with the fact that both of these emotions have high associations with the pleasantness appraisal.
@@ 10 Discussion
A key distinction of our work is that we grounded our experiments on psychological theory and applied MI analysis on in-the-wild examples, rather than relying on synthetically generated simplistic structures, as seen in prior studies Merullo et al. (2024). For example, Wang et al. (2023a) study the Indirect Object Identification task, by considering a fixed input structure, such as “person1 and person2 had fun at school. person2 gave a ring to” where the model is expected to predict “person1”. Hanna et al. (2023) study the “Greater-than" task using a dataset of examples like “The war lasted from 1517 to 15”, where the model is expected to predict any two-digit number larger than 17. While such notable efforts have demonstrated the feasibility of identifying end-to-end specialized circuits for these tasks, effectively interpreting them still requires structured inputs, making it challenging to generalize findings to more naturalistic settings.
Our work also highlights new opportunities for future research. Despite significant advancements in understanding human emotions, debates persist regarding the definition of emotion, the role of cognition in emotion, and the mechanisms underlying emotion inference Ortony et al. (2022); Ellsworth and Scherer (2003); Moors (2013); Barrett (2017). In parallel, cognitive neuroscience has explored the neural basis of emotion in support of differing theoretical perspectives Kragel et al. (2024). The study of LLMs, combined with insights from emotion theory and neuroscience, opens a unique intersection for advancing our understanding of emotions Sievers and Thornton (2024).
@@ 11 Limitations
In this study, we build upon the linear representation hypothesis Mikolov et al. (2013a, b); Levy and Goldberg (2014); Elhage et al. (2022)—the idea that high-level concepts are encoded linearly within model representations Park et al. (2024). This hypothesis is particularly appealing because, if true, it could enable simple and effective methods for interpreting and controlling LLMs—an approach we leveraged to localize and manipulate latent emotion representations. However, despite notable efforts to formalize the notions of linearity Park et al. (2024) and orthogonality Jiang et al. (2024b) in model representations, recent research suggests that not all features are encoded linearly Engels et al. (2025). Further investigation is needed to improve clarity and robustness in this area.
@@ 12 Ethical Impact Statement
This study re-analyzes previously collected, de-identified data that had already undergone ethical review. The dataset is used for investigating the inner mechanisms by which auto-regressive LLMs process emotion. However, caution must be exercised when generalizing these results to models not examined in this work, to superficially similar tasks, or to different languages.
Our analysis highlights potential concerns for those deploying LLMs in high-stakes affective domains or for generating emotionally charged content. Given the risks associated with emotional manipulation by LLMs, it is crucial to develop a deeper understanding of how these models process emotions. To this end, we advocate for further research in this domain to ensure that LLMs align with ethical standards and human-centered AI principles.
Acknowledgments
This work is, in part, supported by the Army Research Office under Cooperative Agreement Number W911NF-25-2-0040. Only staff at ICT were sponsored directly by the Army Research Office. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Office or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.
@@ A.1 Dataset Details
For this project, we employed the crowd-enVENT dataset. Troiano et al. (2023) developed the dataset by asking crowdsource writers to share an event that made them feel a particular emotion. The vignettes varied in length, ranging from short sentences to longer narratives, but the language was predominantly everyday English. The participants were then asked to evaluate their subjective experiences during that event, including their perceived appraisals. Both were rated on a 5-point Likert scale with 5 representing the highest agreement.
@@ A.3 Task Details
For the rest of our analysis, we only focus on the correctly classified examples, which ensures at least 2,700 data points or more across different model architectures. We apply this filtering method to ensure that the samples selected for experimentation are ones where the LLM understood the emotion labeling task, allowing us to properly investigate the underlying mechanisms that led to the models’ selected emotion label. However, we acknowledge that some emotion classes have a limited number of examples after filtering, which may present constraints in our experiments. A preliminary study targeting the bias of such filtering is provided in Appendix C.6, while deeper investigations are left for future studies.
Most experiments in this paper did not have a stochastic nature or behavior, and we used all the available data. The only source of stochasticity in our experiments comes from the train/test split in probe training. However, after extensive experiments, we noticed that the variance for these experiments is extremely low. More precisely, we observe no more than 1.5% STD in classification accuracy across 5 different seeds of each individual probe across all layers, with $R^{2}$ differing by no more than 0.02 between runs for regression tasks. The reported values in the paper figures are therefore only the mean values. Note that we used 5-fold cross-validation for probe hyperparameter tuning; therefore, no extra validation data split was necessary.
@@ C.5 Robustness to Prompt Design
C.6 Investigating the Incorrectly Classified Samples
As mentioned earlier, we only used the dataset instances on which the LLM could predict the self-reported emotion label with no mistakes. We initially adopted this approach to reduce noise in the dataset, given the subjective nature of emotion labeling. Emotion classification can become a very subjective task in some contexts. For example, in many cases, it is hard to distinguish anger from guilt, given the short description provided by the participants. Please see Figure 9, where we show how closely related emotions might be confused. We emphasize that focusing only on correctly classified examples may introduce bias, particularly a bias toward unambiguous emotional content. This is common in much of the prior work in mechanistic interpretability, which focuses on tasks that LLMs perform well on to ensure reliable interpretation.
That being said, analyzing misclassified examples could offer a deeper understanding of model mechanisms. In this section, we conduct a preliminary analysis comparing appraisal patterns between correctly and incorrectly classified cases for each emotion. Interestingly, we found that LLMs tend to exhibit distinct appraisal profiles in misclassified cases, suggesting shifts in the underlying reasoning process.
Table 3 summarizes these findings for three emotions that showed high confusion, presenting the mean and standard deviation of key appraisal dimensions using the appraisal probes we trained, along with t-test comparisons. These results suggest that misclassifications are not simply random errors but may stem from meaningful shifts in how the model evaluates appraisal cues.
Emotion
Pleasantness
(Correct / Miss)
Others agency
(Correct / Miss)
Sadness $1.3\pm 0.5/1.6\pm 0.7$ $2.2\pm 0.7/3.2\pm 0.9$
Joy $4.5\pm 0.5/4.2\pm 0.7$ $3.5\pm 0.7/3.1\pm 0.8$
Guilt $1.9\pm 0.6/1.6\pm 0.8$ $2.8\pm 0.7/3.3\pm 0.7$
Table 3: Comparing the average appraisal values the LLM perceived for correctly classified vs misclassified samples with different emotion labels. All values reported here have p-value less than $0.001$
@@ Appendix F Further Details on Appraisal Modulation
On the other hand, We observe a remarkable alignment with theoretical and intuitive expectations in the distribution shifts associated with interventions on middle layers (specifically layers 9-11). For instance, we observe that increasing the pleasantness appraisal promotes both joy and pride, aligning with the fact that both of these emotions have high associations with the appraisal. Also see Figure 35 for a different visualization.
@@
Finally, we will conclude this section by providing a control experiment in which, instead of using appraisal vectors, we intervene in the activation using a random vector. The results of this experiment are provided in Figure 36, and we see that no psychologically valid pattern is observable. Overall, these findings provide strong evidence that the mid-layers in Llama 3.2 1B meaningfully and directly contribute to cognitive processes related to emotions.
@@
Figure 36: Results of the control experiment in which we randomly sample a vector and add it to the hidden state of the Llama 3.2 1B model at different layers. Each Row shows a different random seed.
@@ Appendix G Code and Compute Resources
Our experiments are conducted using GPU-accelerated compute resources, with hardware such as NVIDIA A100 GPUs. For larger models, our studies are feasible on GPUs with at least 40GB of VRAM, with the full experiment running in approximately 24 hours. For smaller models, GPUs with 12GB of VRAM are sufficient to carry out our analyses efficiently.